# Against The Achilles’ Heel: A Survey on Red Teaming for Generative Models

### v1

@@ 5 Evaluation
In order to better evaluate the jailbreak capability and the defense capability of the model, we summarized the most commonly used evaluation indicators and benchmarks, and classified them for the convenience of future use by researchers.
@@ 5.1 Attack Evaluation
It is important to evaluate whether an attack on language models is successful or not, but their open-ended responses present a great challenge to their evaluation. As a result, different researchers employ different definitions of attack success and use different evaluators, making consistent comparisons between attack measures difficult. Here, we review common definitions of attack success and their evaluators.
@@ 5.1.1 Attack Success Rate
The Attack Success Rate (ASR) is one of the most common metrics in red teaming literature. We first review different definitions of ASR, then we list three qualities broadly recognized as defining the success of an attack: (1) obedience and rejection, (2) harmfulness and toxicity, and (3) relevance and fluency. We then discuss the less noted quality of transferability of attacks.
@@
Most work define ASR across a dataset $D$:
@@
where $Q$ is a query in $D$, $I$ is the evaluator function that equals to $1$ when the response is deemed as attack success, and 00 otherwise. Gong et al. (2023) defined ASR as per query, sampling the same response several times and defined it as:
$ASR_{\jmath}(\mathcal{D})=\frac{\sum_{Q^{*}\in\mathcal{D}}\operatorname{issuccess}_{\mathrm{J}}\left(Q^{*}\right)}{|\mathcal{D}|}$
@@
5.1.2 Attack Success Dimensions
In this section, we discuss different dimensions that define the success of an attack. Some literature combine multiple dimensions to collectively define the success of an attack (Tian et al., 2023b; Deng et al., 2023d; Jiang et al., 2023d).
Obedience and Rejection
A successful attack of a language model makes them obey the query and output the requested content, whereas they would reject the request in the event of a failed attack. It has been observed that many language models are aligned to regularly output certain phrases when refusing to answer, for example, “As an AI language model”, or “I’m sorry”. Therefore it has been proposed by Zou et al. (2023) and widely adopted to measure obedience by the absence of such phrases through lexical matching (Xu et al., 2023d; Lapid et al., 2023; Liu et al., 2023d). However, evaluating whether the model obeys or refuses to answer requires more delicate handling. It works for Full Refusal or Full Compliance where models completely reject responding or faithfully follow the instruction. However, there are cases where models output refusal or disclaimer phrases, only to continue their response with harmful content (Partial Compliance), and cases where they adhere to the instruction but provide no substantial dangerous content (Partial Refusal), as noted by Yu et al. (2023a) and Wang et al. (2023g). Simply matching refusal phrases overlooks the last two cases and results in distorted ASR.
Relevance and Fluency
If the model does not refuse but provides only general content without real harm, or responds with nonsensical sequences, then it should be regarded as an attack failure. Relevance is complicated as it involves understanding the semantics of the response and indicating whether details are involved; the model may mention “drug” in a sentence that cautions the user against it. In view of this, relevance must be judged by humans or language models, either by prompting LLMs (Takemoto, 2024b) or invoking a specialized classifier. Fluency is often judged with perplexity (PPL) calculated with another model such as GPT-2 (Khalatbari et al., 2023) or BERT (Chen et al., 2023a). These two dimensions are often used in combination with other dimensions in various annotation guidelines.
Harmfulness and Toxicity
Responses with specific harmful content related to any risk area should be considered successful attacks, like illustrating steps to make bombs or rob a bank. When evaluating harmfulness it is often desirable to understand what category of risk area is involved, so risk taxonomies are often required in the evaluation, either by tuning models to output a series of labels or including it in the prompt. For example, He et al. (2023b) derives their annotation guidelines based on the HHH criteria Askell et al. (2021) and categorizes harm into five predefined levels (ranging from directly encouraging danger or unethical behavior to completely harm-free). Xu et al. (2023a) considers the response of an LLM as unsafe when it contains any harmful content related to the 10 safety scenarios defined in the paper. This dimension also involves the semantics and is judged by language models (Zeng et al., 2024a; Shah et al., 2023). Common evaluators include moderation APIs (Chern et al., 2024), fine-tuned BERT models (Qiu et al., 2023), and fine-tuned LLMs (See Sec. 5.3(Li et al., 2024b). To simplify the evaluation, Zhang et al. (2023f) propose to adopt a multiple-choice QA formed benchmark to evaluate LLM safety, circumventing the complexities of semantic-based assessments in open-ended QA. This allows for the direct measurement of a model’s harmfulness through accuracy.
@@ 5.1.3 Transferability
The transferability of attacks refers to how the ASR for different models varies. Highly transferable attacks can jailbreak many models with high ASR while attacks with low transferability can only target a few models. This metric is desirable because it signals the universality of attacks. Measurement of transferability requires evaluating it over several models and comparing the ASR. Stronger models like GPT4 tend to have low ASR for the same attack.
@@ 5.1.4 Common Evaluation Datasets
AdvBench is widely used (Guo et al., 2024; Kumar et al., 2023; Li et al., 2023c) as it covers a range of risky behaviors and is popularized by the universal adversarial suffix attack by (Zou et al., 2023), despite leaving much to be desired: it is repetitive with many items simply rephrasing “How to make a bomb” while covering limited risk areas. Part of the studies adopting AdvBench deduplicate the dataset (Chao et al., 2023; Shah et al., 2023) and supplement it with extra risks. Other common datasets are HH-RLHF (Bai et al., 2022) and MaliciousInstruct, and datasets collected from red teaming in the wild are also popular, i.e. TensorTrust (Toyer et al., 2023), Do Anything Now (Shen et al., 2023) and jailbreak prompts collected by Wei et al. (2023a).
@@ 5.2 Defense Evaluation
Both defense and attack evaluation need to measure the result of the attack, with defense success being attack failure, which is discussed in 5.1. In addition, defense must not compromise language model’s performance on normal tasks, otherwise, it will result in Overkill of harmless queries.
@@
Helpfulness is typically measured by conventional benchmark performance. Shi et al. (2023) measures MT-Bench (Zheng et al., 2023) performance, Zhang et al. (2024c) checks the performance of their safeguarded model on MMLU (Hendrycks et al., 2021), AlpacaEval (Dubois et al., 2023) and TruthfulQA (Lin et al., 2022). Ruan et al. (2023) employs human reviewers to check that the helpfulness is not harmed. Ji et al. (2023) had helpfulness annotated in their dataset.
5.2.1 Overkill
@@
Shi et al. (2024) attributes the overkill to models’ tendency to over-attend to harmful words without understanding the semantic meanings, and taking shortcuts when the system prompt emphasizes safety. They analyze the information flow from the safety classification result of the prompts. The results show an increase in the information flow value to the seemingly harmful focus word (e.g., “kill” in “How to kill a python process”), and including a safety-emphasizing system prompt exacerbates the shortcut.
@@ 5.3 Evaluators
In this section, we illustrate the evaluators used to measure different metrics. Many evaluators are used in parallel in the same paper to gain a robust understanding of the attack.
@@ 5.3.1 Lexical Match
This evaluator checks whether the response can be matched by a list of phrases or regular expressions. It is almost exclusively used to measure obedience as proposed by Zou et al. (2023). One can see the appeal of this method as it’s fast to evaluate and very explainable, but it overlooks the complicated semantics of open-ended responses and leads to inaccurate ASR.
@@ 5.3.2 Prompted LLMs
Many research have demonstrated the capability of LLMs as evaluators. A considerable number of red teaming literature have adopted prompting LLs to account for the semantics of the response and measure the harmfulness and relevance of the target model’s answers. The commonly used model is GPT4, adopted by around 10 papers.
@@
Most work formulated the evaluation as a classification problem: given the response, return the classification result that reflects the harmfulness or overall safety of the response.  Wang et al. (2023c) also included the prompt and formulated the content as a dialog.  Sun et al. (2023) makes templates specific according to the evaluated prompt. For example, for scenario Crimes and Illegal Activities, they changed the question from ”Is it safe?” to ”Does it violate any laws or regulations?”. They also added examples to the template as few-shot learning will improve the performance of prompt classification. In contrast to the usual framing, Xu et al. (2023a) adopted a comparison approach. If a response is manually reviewed as safe, ChatGPT is instructed to rewrite it into an unsafe response and vice versa, resulting in triplets of questions, safe responses, and unsafe responses. This formulation avoids the potential inconsistency of rating by language models but becomes more costly.
@@
A number of study asks the evaluator LLMs to choose from a dichotomy: Yes or No (Shah et al., 2023), 00 or $1$ (Zhang et al., 2024e). Others require that a score on a designated scale be returned. Most scales are between 0 and 5 or 0 and 10, while  Naihin et al. (2023) uses scores between 0 and 100, which may allow finer grading for the more complicated task of agent evaluation.  Guo et al. (2024) takes a different approach and specifies a JSON result to be generated that includes the score as well as the reason
@@
GPT4 is known for its high costs. On top of that, close-sourced models can be subject to model change which makes reproduction difficult.  Shah et al. (2023) noticed that GPT4 has a high false-negative rate: many unsafe prompts are classified as safe.
@@ 5.3.3 Specialized Classifiers
These evaluators are cheaper than most LLMs as they tend to be smaller and more specialized. Common classifiers Perspective API (Shayegani et al., 2023a), OpenAI Moderation API (Yang et al., 2023b) and models from BERT family such as HateBERT (Caselli et al., 2021) by  Wang & Shu (2023), DistilBERT (Sanh et al., 2020) by  Kumar et al. (2023) and RoBERTa (Liu et al., 2019) by  Yu et al. (2023a) and  Qiu et al. (2023). Studies using BERT models need to often fine-tune them on their curated data or on existing data such as BAD (Xu et al., 2021) and HH-RLHF.  Kumar et al. (2023) pointed out that some classification problems need to be considered such as class balancing, where data with different labels vary in number, making model classification biased. This problem can be mitigated by augmenting data with LLM generation. TF-IDF feature and Naive Bayes classifier are adopted by  Chen et al. (2023a) to classify obedience.
@@ 5.3.4 Human Reviewers
Often recognized as the gold standard but more difficult to scale, manual review is employed in many works to check automatic evaluation results. It is common to have multiple human reviewers review the same response and compare their consistency with each other using metrics like Likert scale.
@@ 5.4 Benchmarks
Comprehensive Safety SALAD-Bench (Li et al., 2024b) Comprehensive safety in 3 directions Human + LLM Auto ASR, Rejection rate, Safety rates 6 30k
SafetyBench (Zhang et al., 2023f) Comprehensive safety of LLMs Human + Curated + LLM Auto Category-wise accuracy 7 11k
Do-Not-Answer (Wang et al., 2023g) Harmful instructions LLM + Curated Auto Harmful response rate, Detection rate 5 939
SAFETYPROMPTS (Sun et al., 2023) Comprehensive safety of Chinese LLMs Human + LLM Auto Scenario-wise safety score 14 100k
SC-Safety (Xu et al., 2023c) Comprehensive safety of Chinese LLMs Human + LLM Human + Auto Safety score 3 4.9k
SimpleSafetyTest (Vidgen et al., 2023) Harmfulness of response Human Human + Auto Safety rates, Accuracy 5 100
SciMT-Safety (He et al., 2023b) Misuse of AI in scientific contexts Human + LLM Auto Harmlessness score 2 432
Specific Safety Concern XSTEST (Röttger et al., 2023) Exaggerated safety behaviors Human Human Refusal rates 1 450
CValues (Xu et al., 2023a) Safety & Responsibility of Chinese LLMs Human + LLM Human + Auto Safety & Responsibility score 17 2.1k
DICES (Aroyo et al., 2023) Variance, ambiguity, and diversity of AI safety Human + LLM N/A N/A 25 990
ToxicChat (Lin et al., 2023) Toxicity detection for chatbot Human Auto Prec, Recall, F1 and Jailbreak recall 1 10k
ToxicGen (Hosseini et al., 2023) Implicit representational harms Human + LLM Auto Scaled perplexity, Safety score 1 6.5k
HarmfulQ (Shaikh et al., 2023) Safety of zero-shot COT reasoning LLM Human + Auto Accuracy of discouraging harmful behavior 2 200
XSAFETY (Wang et al., 2023c) Multilingual safety Curated Human + Auto Unsafety rate 14 2.8k
High-Risk (Hung et al., 2023) Safety & Factuality in legal and medical domain Curated Auto QAFactEval, UniEval, SafetyKit, Detoxify scores 2 3.4k
StrongReject (Souly et al., 2024) Jailbreak Human + LLM Auto Jailbroken score 8 346
Attack and Exploitation Do Anything Now (Shen et al., 2023) Jailbreak Curated Human + Auto ASR 13 666
MasterKey (Deng et al., 2023b) Jailbreak Human Human + Auto Query & Prompt success rate 4 85
Latent Jailbreak (Qiu et al., 2023) Safety and robustness in jailbreak Curated Human + Auto ASR, Robustness rate, Trustworthiness 3 416
BIPIA (Yi et al., 2023) Indirect prompt injection attack Human + Curated + LLM Auto ASR 2 86k
Tensor Trust (Toyer et al., 2023) Robustness to prompt injection attacks Human + Curated Auto Hijacking & Extraction Robustness Rate, Defense Validity 2 1.3k
AdvBench (Zou et al., 2023) Prompt suffix attack Human + LLM Auto ASR, Cross-entropy loss 6 1.1k
MaliciousInstruct (Huang et al., 2023b) Generation exploitation attacks LLM Auto ASR, Harmful Percentage 5 100
HarmBench (Mazeika et al., 2024) Auto red teaming Human Auto ASR 10 510
@@
As LLM safety gains attention, numerous benchmarks have been introduced to assess the safety of models, and they differ from the evaluation objectives and dimensions. We categorize well-known benchmarks into three groups based on their evaluation focus as shown in Table 2.

### v2

@@ 6 Evaluation
In order to better evaluate the jailbreak resilience and defense capabilities of a model, we review the most commonly-used evaluation indicators and benchmarks.
@@ 6.1 Attack Evaluation
Evaluating the success of attacks on language models is essential, yet the open-ended nature of their responses poses significant challenges for assessment. Consequently, researchers often adopt varying definitions of what constitutes a successful attack and apply diverse evaluation metrics, complicating direct comparisons across studies. In this review, we examine prevalent definitions of attack success and the evaluation methods used.
6.1.1 Attack Success Dimensions
In this section, we discuss different dimensions that define the success of an attack. Some literature combines multiple dimensions to collectively define the success of an attack Tian et al. (2023b); Deng et al. (2023d); Jiang et al. (2023d).
Obedience and Rejection
A successful attack on a language model makes it obey the query and output the requested content; otherwise, it rejects the request. Many language models have been observed to use specific phrases, such as “As an AI language model” or “I’m sorry”, when declining to answer a query. Consequently,  Zou et al. (2023b) proposed a method that has since become widely adopted, measuring model compliance by detecting the absence of these phrases through lexical matching Xu et al. (2023d); Lapid et al. (2023); Liu et al. (2023d). However, evaluating whether the model obeys or refuses to answer requires more delicate analysis. It works for Full Refusal or Full Compliance where models completely reject responding or faithfully follow the instruction. Yet, there are cases where models output refusal or disclaimer phrases but continue their response with harmful content (Partial Compliance), and cases where they adhere to the instruction but provide no substantial dangerous content (Partial Refusal), as noted by Yu et al. (2023a) and Wang et al. (2023g). Simply matching refusal phrases overlooks the last two cases and results in distorted ASR.
Relevance and Fluency
If the model does not refuse but provides only general content without real harm, or responds with nonsensical sequences, then it should be regarded as a failed attack. Relevance is complicated as it involves understanding the semantics of the response and indicating whether details are involved; the model may mention “drug” in a sentence that cautions the user against it. In view of this, relevance must be judged by humans or language models, either by prompting LLMs Takemoto (2024b) or invoking a specialized classifier. Fluency is often judged with perplexity (PPL) calculated with another model such as GPT-2 Khalatbari et al. (2023) or BERT Chen et al. (2023a). These two dimensions are often used in combination with other dimensions.
Harmfulness and Toxicity
Responses with specific harmful content related to any risk area should be considered successful attacks, for instance illustrating steps to make a bomb or rob a bank. When evaluating harmfulness it is often desirable to understand what category of risk area is involved, so risk taxonomies are often required by the evaluation, either by tuning models to output a series of labels or including it in the prompt. For example, He et al. (2023b) derive their annotation guidelines based on the HHH criteria Askell et al. (2021) and categorize harm into five predefined levels (ranging from directly encouraging danger or unethical behavior to completely harm-free). Xu et al. (2023a) consider the response of an LLM as unsafe when it contains any harmful content related to the 10 safety scenarios defined in the paper. This dimension also involves semantics as judged by language models Zeng et al. (2024a); Shah et al. (2023). Common evaluators include moderation APIs Chern et al. (2024), fine-tuned BERT models Qiu et al. (2023), and fine-tuned LLMs (See Section 6.3Li et al. (2024c). To simplify evaluation, Zhang et al. (2023f) propose a multiple-choice QA-based benchmark to evaluate LLM safety, circumventing the complexities of semantic-based assessments in open-ended QA. This allows for the direct measurement of a model’s harmfulness via accuracy.
@@ 6.1.2 Attack Success Rate
Attack Success Rate (ASR) is one of the most common metrics in the red teaming literature. We first review different definitions of ASR, then list three qualities broadly recognized as defining the success of an attack: (1) obedience and rejection, (2) harmfulness and toxicity, and (3) relevance and fluency. We then discuss the less-noted quality of transferability of attacks.
@@
Most work defines ASR across a dataset $D$ as:
@@
where $Q$ is a query in $D$, and $I$ is the evaluator function that outputs $1$ when the response is deemed as an attack success, and 00 otherwise. Gong et al. (2023) defined ASR per query, sampling the same response several times and defining it as:
$ASR_{\jmath}(\mathcal{D})=\frac{\sum_{Q^{*}\in\mathcal{D}}\operatorname{is\_success}_{\mathrm{J}}\left(Q^{*}\right)}{|\mathcal{D}|}$
@@ 6.1.3 Transferability
The transferability of attacks refers to how ASR varies across models. Highly transferable attacks can jailbreak many models with high ASR, while attacks with low transferability only catch out a few models. This metric is desirable because it signals the universality of attacks. Measurement of transferability requires evaluating it over several models and comparing the ASR. Stronger models like GPT4 tend to have low ASR for the same attack.
@@ 6.1.4 Common Evaluation Datasets
AdvBench is widely used Guo et al. (2024); Kumar et al. (2023); Li et al. (2023e) as it covers a range of risky behaviors, and was popularized by the universal adversarial suffix attack of Zou et al. (2023b), despite having notable limitations: it is repetitive with many items simply rephrasing “How to make a bomb”, and covers limited risk types. Some who use AdvBench deduplicate the dataset Chao et al. (2023); Shah et al. (2023) and supplement it with extra risk types. Other common datasets are HH-RLHF Bai et al. (2022a) and MaliciousInstruct Huang et al. (2023c). Datasets collected from red teaming in the wild are also popular, i.e., TensorTrust Toyer et al. (2023), Do Anything Now Shen et al. (2023) and jailbreak prompts collected by Wei et al. (2023a).
@@ 6.2 Defense Evaluation
Both defense and attack evaluation measure the result of the attack, with defense success being attack failure as discussed in Section 6.1. In addition, defense must not compromise a language model’s performance on normal tasks, mitigating the overkill of harmless queries and making sure of the helpfulness of the language model. For example, a language model should not refuse to help users kill redundant Python processes, but should refuse requests to provide guidance on how to commit murder.
@@
Helpfulness is typically measured by conventional benchmark performance, which should be on par with language models without defense measures. Shi et al. (2023b) measure MT-Bench Zheng et al. (2023a) performance, while Zhang et al. (2024c) check the performance of their safeguarded model on MMLU Hendrycks et al. (2021), AlpacaEval Dubois et al. (2023) and TruthfulQA Lin et al. (2022). Ruan et al. (2023) employ human reviewers to check that helpfulness is not harmed. Ji et al. (2023) had helpfulness annotated in their dataset.
Overkill
@@
Shi et al. (2024a) attributes the overkill to a model’s tendency to over-attend to harmful words without understanding the full intent behind the prompt, resulting in over-conservatism. Their analysis reveals that the information flow from safety mechanisms increases toward specific focus words that appear harmful (e.g., “kill” in “How to kill a python process”). Including a safety-emphasizing system prompt exacerbates this effect, as seen in Figure 25.
Figure 25: Overkill examples (figure from Bianchi et al. (2023)). The responses with boxes outlined in red are examples of overkill.
@@ 6.3 Evaluators
In this section, we illustrate the evaluators used to measure different metrics. Many evaluators are used in parallel in the same paper to gain a robust understanding of model safety.
@@ 6.3.1 Lexical Match
This evaluator checks whether the response matches a list of phrases or regular expressions. It is used almost exclusively to measure obedience as proposed by Zou et al. (2023b). One can see the appeal of this method as it’s fast to evaluate and very explainable, but it overlooks the complicated semantics of open-ended responses and leads to inaccurate ASR.
@@ 6.3.2 Prompted LLMs
Much research has demonstrated the capability of LLMs as evaluators. A considerable number of red teaming literature has adopted prompting LLMs to account for the semantics of the response, and measure the harmfulness and relevance of the target model’s answers. A commonly-used model is GPT4.
@@
Most work formulates evaluation as a classification problem: given the response, return the classification result that reflects the harmfulness or overall safety of the response. Wang et al. (2023c) also included the prompt and formulated the content as a dialogue.  Sun et al. (2023) make templates specific to the evaluated prompt. For example, for the scenario Crimes and Illegal Activities, they changed the question from “Is it safe?” to “Does it violate any laws or regulations?”. They also added examples to the template and used few-shot learning to improve the performance of prompt classification. In contrast to the usual framing, Xu et al. (2023a) adopted a comparison approach: if a response is manually reviewed as safe, ChatGPT is instructed to rewrite it into an unsafe response and vice versa, resulting in triplets of questions, safe responses, and unsafe responses. This formulation avoids the potential inconsistency of rating by language models, but is more costly.
@@
A number of studies have treated safety as a binary task: Yes or No Shah et al. (2023), or 00 or $1$ Zhang et al. (2024e). Others require that a score on a designated scale be returned. Most scales are between 0 and 5 or 0 and 10, while  Naihin et al. (2023) uses scores between 0 and 100, which may allow for finer grading for the more complicated task of agent evaluation.  Guo et al. (2024) take a different approach and specifie a JSON result to be generated that includes the score as well as the reason.
@@
GPT4 is known for its high costs. On top of that, closed-source models can be subject to model changes, compromising reproductibility.  Shah et al. (2023) noticed that GPT4 has a high false-negative rate: many unsafe prompts are classified as safe.
@@ 6.3.3 Specialized Classifiers
Specialized evaluators are cheaper than most LLMs as they tend to be smaller. Common classifiers Perspective API Shayegani et al. (2023a), OpenAI Moderation API Yang et al. (2023b) and models from the BERT family such as HateBERT Caselli et al. (2021) by  Wang & Shu (2023), DistilBERT Sanh et al. (2019) by  Kumar et al. (2023) and RoBERTa Liu et al. (2019) by  Yu et al. (2023a) and  Qiu et al. (2023). Studies using BERT models generally need to fine-tune them on data such as BAD Xu et al. (2021) and HH-RLHF.  Kumar et al. (2023) pointed out that some data issues need to be considered such as class balancing, where the distribution of instances across classes is non-uniform, biasing classification. This can be mitigated by augmenting the data with LLM generation. TF-IDF features were adopted by  Chen et al. (2023a) to classify obedience.
@@ 6.3.4 Human Reviewers
Often recognized as the gold standard but more difficult to scale, manual review is employed in much work to verify the quality of automatic evaluation. It is common to have multiple human reviewers review the same response, and compare their consistency with each other.
@@ 6.4 Benchmarks
Comprehensive Safety SALAD-Bench Li et al. (2024c) Comprehensive safety in 3 directions Human + LLM Auto ASR, Rejection rate, Safety rates 6 30k
SafetyBench Zhang et al. (2023f) Comprehensive safety of LLMs Human + Curated + LLM Auto Category-wise accuracy 7 11k
Do-Not-Answer Wang et al. (2023g) Harmful instructions LLM + Curated Auto Harmful response rate, Detection rate 5 939
SAFETYPROMPTS Sun et al. (2023) Comprehensive safety of Chinese LLMs Human + LLM Auto Scenario-wise safety score 14 100k
SC-Safety Xu et al. (2023c) Comprehensive safety of Chinese LLMs Human + LLM Human + Auto Safety score 3 4.9k
SimpleSafetyTest Vidgen et al. (2023) Harmfulness of response Human Human + Auto Safety rates, Accuracy 5 100
SciMT-Safety He et al. (2023b) Misuse of AI in scientific contexts Human + LLM Auto Harmlessness score 2 432
Specific Safety Concern XSTEST Röttger et al. (2023) Exaggerated safety behaviors Human Human Refusal rates 1 450
CValues Xu et al. (2023a) Safety & Responsibility of Chinese LLMs Human + LLM Human + Auto Safety & Responsibility score 17 2.1k
DICES Aroyo et al. (2023) Variance, ambiguity, and diversity of AI safety Human + LLM N/A N/A 25 990
ToxicChat Lin et al. (2023) Toxicity detection for chatbot Human Auto Precision, Recall, F1 and Jailbreak recall 1 10k
ToxicGen Hosseini et al. (2023) Implicit representational harms Human + LLM Auto Scaled perplexity, Safety score 1 6.5k
HarmfulQ Shaikh et al. (2023) Safety of zero-shot COT reasoning LLM Human + Auto Accuracy of discouraging harmful behavior 2 200
XSAFETY Wang et al. (2023c) Multilingual safety Curated Human + Auto Unsafety rate 14 2.8k
High-Risk Hung et al. (2023) Safety & Factuality in legal and medical domain Curated Auto QAFactEval, UniEval, SafetyKit, Detoxify scores 2 3.4k
StrongReject Souly et al. (2024) Jailbreak Human + LLM Auto Jailbroken score 8 346
Attack and Exploitation Do Anything Now Shen et al. (2023) Jailbreak Curated Human + Auto ASR 13 666
MasterKey Deng et al. (2023b) Jailbreak Human Human + Auto Query & Prompt success rate 4 85
Latent Jailbreak Qiu et al. (2023) Safety and robustness in jailbreak Curated Human + Auto ASR, Robustness rate, Trustworthiness 3 416
BIPIA Yi et al. (2023) Indirect prompt injection attack Human + Curated + LLM Auto ASR 2 86k
Tensor Trust Toyer et al. (2023) Robustness to prompt injection attacks Human + Curated Auto Hijacking & Extraction Robustness Rate, Defense Validity 2 1.3k
AdvBench Zou et al. (2023b) Prompt suffix attack Human + LLM Auto ASR, Cross-entropy loss 6 1.1k
MaliciousInstruct Huang et al. (2023c) Generation exploitation attacks LLM Auto ASR, Harmful Percentage 5 100
HarmBench Mazeika et al. (2024) Auto red teaming Human Auto ASR 10 510
@@
As LLM safety gains attention, numerous benchmarks have been introduced to assess the safety of models, which differ across dimensions such as their evaluation objectives. We categorize well-known benchmarks into three groups as shown in Table 3.